# Empirical Analysis of Pull Requests for Google Summer of Code

### v1

@@ VI Discussion and Implications for Software Engineering Education
The results of the analysis on GitHub pull requests and review comments can have significant implications for software engineering education. The tasks and feedback topics identified in review comments can aid educators to better prepare students for real-world software development practices. The findings highlight key areas that should be emphasized in software engineering curricula, as outlined below.
@@
    Importance of code reviews in collaborative software development. The study reveals the importance of peer reviews in ensuring the quality of code contributions. For software engineering education, this means integrating code reviews into course projects and emphasizing communication skills.
@@
    Code organization and structure. The findings reveal that reviewers frequently comment on the structure and organization of code, such as where files should be placed and how modules should be organized. This highlights the need to emphasize code organization in software engineering education by 1) teaching modular design and file organization, and 2) following real-world practices in project layout.
@@
    Testing and test-driven development The study shows that a significant portion of review feedback is related to missing or inadequate testing. This suggests that educators should place a stronger emphasis on teaching comprehensive testing practices and emphasizing edge cases.
@@
    Error handling and robustness. Feedback in the review comments frequently addresses issues with error handling, suggesting that students often underestimate its importance. In response, educators should teach best practices for error handling and include error handling in assessment criteria.
@@
    Performance optimization. Reviewers also suggest performance optimizations or improvement on the coding logic. For software engineering education, this implies teaching algorithmic efficiency and introducing profiling tools.
@@
    Refactoring and continuous improvement. Reviewers frequently suggest refactoring code for better readability, efficiency, and maintainability. In education, this highlights the need to encourage refactoring as a habit and teach clean code principles.
@@
    Prepare Students for open-source contributions. Many of the findings reflect practices commonly seen in open-source development environments, such as GitHub. The alignment of software engineering education with these practices can prepare students better to contribute to open-source projects. Educators can use real-world GitHub projects as learning tools, and incorporate the use of version control for collaboration in software engineering classes.
@@
    Learning to give and receive feedback. The study shows that review comments often provide specific, constructive feedback on how to improve code. Software engineering education should emphasize 1) ability of students to provide constructive feedback and 2) receiving and acting on feedback provided by peers.
The findings of this study offer several key insights for software engineering education. When educators focus on real-world practices with respect to code organization, testing, error handling, performance optimization, and collaboration; they can prepare students better for the challenges they will face in the software industry. The integration of code reviews, refactoring exercises, and open-source contributions into the curriculum will likely foster a culture of continuous improvement and collaboration; thereby, equipping future software engineers with the tools they need for success. These educational adjustments will help students internalize best practices in software development, make them better collaborators, more thoughtful developers, and ultimately more successful engineers in both open-source and proprietary software environments.
@@ VII Threats to Validity
There are several potential threats to the validity of our findings. These threats can affect the reliability and generalizability of the findings. The following subsections discuss the potential threats to validity of the results in terms of construct validity, internal validity, external validity, and conclusion validity.
@@ VII-A Construct Validity
  • NLP limitations in identifying themes. The natural language processing (NLP) techniques used in the study, such as keyword extraction, tokenization, and topic modelling may not fully capture the nuances and context of review comments. For instance, comments that use implicit language, technical jargon, or sarcasm might be misinterpreted by simple keyword-based approaches. To minimize this threat, we manually validate a sample of comments. However, improving the NLP model by incorporating more advanced models (e.g., transformer-based models) could improve accuracy in capturing nuanced feedback.
  • Pull request descriptions may not fully capture the nature of changes. Pull request descriptions vary widely in quality and completeness. Some pull request may lack detailed explanations of the changes, leading to incomplete or inaccurate summary of tasks.
@@ VII-B Internal Validity
  • Bias in data collection. The dataset used for the study was sourced from only GitHub repositories of projects that participated in the Google Summer of Code program. Since this program in being run by only Google, there is possibility that the projects follow specific coding practices, use particular technologies, or operate in niche domains. Hence, it is possible that the findings may not be representative of broader software development practices.
@@ VII-C External Validity
  • Generalizability across different projects and domains. The study may not generalize well to all software projects. Different software projects (e.g., open-source vs. proprietary, small vs. large projects, front-end vs. back-end development) have distinct workflows, contributor behaviors, and review cultures. Therefore, findings from a single type of project may not apply broadly.
  • Varying review practices across teams. Review practices may differ based on team norms, the development process (e.g., agile vs. waterfall), or company policies. A highly structured review process may yield different types of feedback than a more informal review process.
  • Changes over time. The feedback practices and development workflows on GitHub may evolve over time, influenced by the introduction of new tools (e.g., code linters, continuous integration pipelines) or changes in the development community’s culture. This temporal aspect might limit the applicability of findings from historical PR data to current or future practices.
@@ VII-D Conclusion Validity
  • Unaccounted confounding factors. Confounding factors such as the complexity of the code changes, the experience level of the contributor, or the urgency of the PR could influence the nature of the feedback provided. If these factors are not accounted for, the conclusions may be skewed.
The study provides valuable insights into pull requests and code review comments. However, the validity of the findings could be affected by the aforementioned threats. We have used a variety of approaches (such as manual validation of some of the comments) to minimize these threats and ensure careful interpretation of the results. In the future we will expand the dataset, refine the analytical techniques, and account for variability in development practices to enhance the validity and generalizability of the findings.
@@ VIII Conclusion
This study provides an in-depth analysis of GitHub pull requests and review comments created during the Google Summer of Code program from 2020 to 2023. The study offers insights into common tasks and feedback themes. The results shows that pull requests often focus on feature additions, bug fixes, and documentation updates. The review comments frequently emphasize the need for improved code organization, additional tests, performance optimizations, and better error handling.
These findings can help developers and teams enhance their review practices by focusing on key areas such as test coverage, performance improvements, and maintaining a well-structured codebase. These results also show that there is need to enhance the software engineering curriculum to include emphasis on code reviews, code organization, testing, and best practices from open source repositories in the curriculum.
In the future ,we plan to expand the dataset to include the available dataset for all the years in the GSoC program. We also plan to study the evolution of the tasks and comments across these years. Finally, we plan to explore other summer programs and hackatons that target students to validate the generalizability of this study findings across different repositories and software domains.

### v2

@@ VI Discussion and Implications for Software Engineering Education
The results of the analysis on GSoC pull requests and review comments have significant implications for software engineering education. The identified tasks, and feedback from review comments can help educators better prepare students for real-world software development practices. The findings highlight key areas that should be emphasized in software engineering curricula, as outlined below.
@@
    Importance of code reviews in collaborative software development. The study underscores the critical role of peer reviews in ensuring the quality of code contributions. For software engineering education, this suggests integrating code reviews into course projects and emphasizing communication skills to facilitate effective collaboration.
@@
    Code organization and structure. The findings show that reviewers frequently comment on the structure and organization of code, such as file placement and module organization. This highlights the need to emphasize code organization in software engineering education by 1) teaching modular design and file organization, and 2) following real-world practices in project layout within software engineering courses.
@@
    Testing and test-driven development The study reveals that a significant portion of review feedback pertains to missing or inadequate testing. Educators should therefore place greater emphasis on teaching comprehensive testing practices, including test-driven development and addressing edge cases.
@@
    Error handling and robustness. Review comments often address issues with error handling, suggesting that students may underestimate its importance. Educators should teach best practices for error handling and include this aspect as part of assessment criteria.
@@
    Performance optimization. Reviewers also suggest performance optimizations or improvements in coding logic. For software engineering education, this highlights the importance of teaching algorithmic efficiency and introducing profiling tools to identify performance bottlenecks.
@@
    Refactoring and continuous improvement. Reviewers often recommend refactoring code for improved readability, efficiency, and maintainability. This underscores the need to encourage refactoring as a regular habit in software engineering education while teaching clean code principles.
@@
    Preparing Students for open-source contributions. Many findings reflect practices commonly observed in open-source development environments, such as GitHub. The alignment of software engineering education with these practices can better prepare students for contributing to open-source projects. Educators can use real-world GitHub projects as teaching tools, and incorporate version control systems to stimulate collaboration in software engineering classes.
@@
    Learning to give and receive feedback. The study highlights the value of specific and constructive feedback in review comments. Software engineering education should emphasize 1) the ability of students to provide constructive feedback and 2) the importance of receiving and acting on feedback from peers.
The findings of this study offer several key insights for software engineering education. When educators focus on real-world practices such as code organization, testing, error handling, performance optimization, and collaboration; they can better prepare students for the challenges of the software industry. The integration of code reviews, refactoring exercises, and open-source contributions into the curriculum will likely foster a culture of continuous improvement and collaboration. This approach equips future software engineers with the skills and mindset they need for success. These educational adjustments will help students internalize best practices in software development, become better collaborators, and emerge as thoughtful, effective developers in both open-source and proprietary software environments.
@@ VII Threats to Validity
There are several potential threats to the validity of our findings. These threats can affect the reliability and generalization of the results. The following subsections discuss the potential threats to the validity of the results in terms of construct validity, internal validity, external validity, and conclusion validity.
@@ VII-A Construct Validity
  • NLP limitations in identifying themes. The natural language processing (NLP) techniques used in the study, such as keyword extraction, tokenization, and topic modeling may not fully capture the nuances and context of review comments. For instance, comments that use implicit language, technical jargon, or sarcasm might be misinterpreted by keyword-based approaches. To minimize this threat, we manually validate a sample of comments. However, incorporating more advanced models (e.g., transformer-based models) could improve the accuracy of capturing nuanced feedback.
  • Pull request descriptions may not fully capture the nature of changes. Pull request descriptions vary widely in quality and completeness. Some pull request may lack detailed explanations of the changes, leading to incomplete or inaccurate summaries of tasks.
@@ VII-B Internal Validity
  • Bias in data collection. The dataset used for the study was sourced solely from GitHub repositories of projects that participated in the GSoC program. Since this program in being run by only Google, there is a possibility that the projects follow specific coding practices, use particular technologies, or operate in niche domains. Hence, the findings may not be representative of broader software development practices.
@@ VII-C External Validity
  • Generalizability across different projects and domains. The study may not generalize well to all software projects. Different types of software projects (e.g., open-source vs. proprietary, small vs. large projects, front-end vs. back-end development) have distinct workflows, contributor behaviors, and review cultures. Therefore, findings from a single type of project may not apply broadly.
  • Varying review practices across teams. Review practices may differ based on team norms, the development process (e.g., agile vs. waterfall), or company policies. A highly structured review process may yield different types of feedback compared to a more informal review process.
  • Changes over time. Feedback practices and development workflows on GitHub may evolve over time, influenced by the introduction of new tools (e.g., code linters, continuous integration pipelines) or changes in the development community’s culture. This temporal aspect might limit the applicability of findings from historical pull request data to current or future practices.
@@ VII-D Conclusion Validity
  • Unaccounted confounding factors. Confounding factors such as the complexity of the code changes, the experience level of the contributor, or the urgency of the pull request could influence the nature of the feedback provided. If these factors are not accounted for, the conclusions may be skewed.
The study provides valuable insights into pull requests and code review comments. However, the validity of the findings could be affected by the aforementioned threats. We have used various approaches (such as manual validation of some of the comments) to minimize these threats and ensure careful interpretation of the results. In the future, we will expand the dataset, refine the analytical techniques, and account for variability in development practices to enhance the validity and generalizability of the findings.
@@ VIII Conclusion
This study provides an in-depth analysis of GitHub pull requests and review comments created during the Google Summer of Code program from 2020 to 2023. The study offers insights into common tasks and feedback themes. The results shows that pull requests frequently focus on feature additions, bug fixes, and documentation updates. The review comments often emphasize the need for improved code organization, additional tests, performance optimizations, and better error handling.
These findings can help developers and teams enhance their review practices by focusing on key areas such as test coverage, performance improvements, and maintaining a well-structured codebase. The results also highlights the need to enhance the software engineering curriculum by incorporating a stronger emphasis on code reviews, code organization, testing, and best practices from open-source repositories.
In the future, we plan to expand the dataset to include data for all available years of the GSoC program. Additionally, we aim to study the evolution of tasks and comments across these years. Finally, we intend to explore other summer programs and hackathons that target students to validate the generalizability of this study’s findings across different repositories and software domains.